Mathematics 132 – Computer Lab Groups

1) Alexandria Andrew and Megan Hyland 2) Luke Barr, Greg Progri, and Gonzalo Vera 3) Amelia Cianci and Michelle DiBlasi

4) Charles Crawley and Jason Tartaglione 5) Isa Dattilo and Gilberto Gomez

6) Caitlin Deitz and Molly O’Keefe

7) Ashley Domingue and Christine Sturdevant 8) Robert Dudley and Jordan Forester

9) Tamara Edme and Kristine Sullivan

10) Erika Furtado and Katrina Muth

11) Shailin Hogan and Markenya Mirander

12) Carissa Lawrence and Loyola Pasiewicz 13) Lauren Nakhoul and Nell Turner

14) Tara Richards and Dominick Sciame

15) Vanessa Dagley and Kristen Oats

16) Aaron Rowden, Mike Thater